# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v2

@@ 5.3 Causal3DIdent Dataset
We conduct our experiments using the Causal3DIdent dataset, a semi-synthetic benchmark widely used in the causal representation learning literature (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023). This dataset enables explicit enforcement of causal structures over latent variables. The generative process for images is governed by a set of 10 latent variables, comprising 3 discrete factors—object shape (shape), and object positions along the horizontal (x_pos) and vertical (y_pos) axes—and 7 continuous factors: object color (color); spotlight position (s_pos) and color (s_color); background color (b_color); and object rotation angles (alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific variables, and the remaining factors as semantic variables, which follow a predefined causal structure illustrated in Figure 14.
@@
Evaluation metrics include $R^{2}$ for continuous latent variables and MCC for discrete latent variables. Further details on experimental setup, including dataset specifications, encoder architectures, and training parameters, are provided in § E.1.
@@ Results.
Figure 6 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably captured across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$. Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure illustrated in Figure 14. Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in § E.2.
@@ 6 Related Work
We now contextualize our work within four key lines of related research: theoretical analysis of contrastive learning, vision–language models, identifiability in latent variable models, and invariant representation learning.
@@ 7 Conclusion and Discussion
In this work, we present a formal analysis of misalignment in MMCL, examining its impact on learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, systematically discarding misaligned latent variables. When image-text pairs exhibit selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights underscore the need for multimodal learning frameworks that either mitigate misalignment or leverage beneficial biases to improve representation learning in real-world settings (see App. G for further discussion).
@@ C.1 Detailed Experimental Setup
We decompose the latent space into a semantic subspace $\mathcal{S}\subseteq\mathbb{R}^{10}$ and two modality-specific latent subspaces, $\mathcal{M}_{x}\subseteq\mathbb{R}^{5}$ and $\mathcal{M}_{t}\subseteq\mathbb{R}^{5}$. Latent variables are sampled from multivariate Gaussian distributions: $\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}})$, $\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}})$, and $\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}})$.
In the independent semantic variables setting, we set $\Sigma_{\mathbf{s}}=I_{10}$, so that $\mathbf{s}$ follows a factorized standard Gaussian distribution, consistent with nonlinear ICA settings (Khemakhem et al., 2020; Sorrenson et al., 2020). In the dependent semantic variables setting, we sample $\Sigma_{\mathbf{s}}$ from a Wishart distribution with an identity scale matrix and degrees of freedom equal to $10$, introducing an unknown causal structure among semantic latent variables, as in prior causal representation learning works (Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023).
For modality-specific variables, distinct covariance matrices $\Sigma_{\mathbf{m}_{x}}$ and $\Sigma_{\mathbf{m}_{t}}$ are independently sampled from a Wishart distribution with an identity scale matrix and degrees of freedom equal to $5$.
@@
Since the semantic subspace has dimension $10$, the range of selection biases $\theta$ is given by $[2^{10}-1]$, resulting in $1023$ possible subsets. To systematically analyze the influence of selection bias, we choose 10 representative scenarios, detailed in Table 2.
For perturbation biases, we fix a full-selection scenario with $\theta=1023$, which also yields $1023$ possible perturbation settings. Similarly, we select 10 representative perturbation-bias scenarios, described in Table 3. Additive Gaussian noise
$\tilde{s}_{i}=s_{i}+\mathcal{N}(0,\Sigma_{\epsilon})$
is applied independently with probability $0.75$ to each semantic dimension $i\in\mathbb{I}_{\rho}$ in modality $\mathbf{t}$, simulating random perturbations. The noise covariance $\Sigma_{\epsilon}$ is sampled from a Wishart distribution with an identity scale matrix and degrees of freedom equal to $10$.
Results from these 20 diverse settings validate our theoretical claims. To explicitly study the joint effect of selection and perturbation biases, we consider the scenario $\theta=968$ for selection bias (excluding the last two indices) and $\rho=12\mid\theta=968$ for perturbation bias (perturbing the first two indices).
@@ Appendix D Experiment Details on MPI3D-Complex Dataset
Figure 11: Random image samples from the MPI3D-Complex dataset. The images are captured by a camera in a controlled environment.
@@ D.1 Detailed Experimental Setup
The MPI3D-Complex dataset (Gondal et al., 2019) contains 460,800 real-world images of resolution $64\times 64\times 3$, captured using a camera in a controlled environment. The dataset spans all possible combinations of seven mutually independent, discrete latent factors, as detailed in Table 6. Representative image samples are shown in Figure 11.
Due to their low-level spatial nature and the difficulty in directly mapping them to semantic descriptions, we treat the horizontal and vertical axes (hori., vert.) as image-specific factors. The remaining factors are considered latent semantic variables for the purposes of our experiments.
@@
Under unbiased conditions, the text semantic variables are identical to the image semantic variables. For biased settings, we include only the content words corresponding to the selected true image semantic variables, as specified in Table 6.
To introduce linguistic variation, we employ three manually designed text templates for text generation. The choice of template is controlled by a text-specific factor $m_{x}$, sampled from a uniform distribution:
$m_{x}\sim\mathrm{Uniform}(\{1,2,3\}).$
@@
In selection bias settings (where the perturbable set $\mathbb{I}_{\rho}$ is empty), textual descriptions include only the content words corresponding to a subset of the true image semantic variables. The specific text generation templates associated with each selection setting are detailed in Table 8.
In perturbation bias settings (where all semantic factors are selected), we introduce misalignment by randomly altering the values of the perturbable semantic variables in the text: each such value is replaced with a randomly sampled value from its support with probability $0.9$.
@@ E.1 Detailed Experimental Setup
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
For both the image and text encoders, we adopt the same architectures used in the MPI3D-Complex experiments. The encoding dimensionality is adjusted according to the bias setting: for selection settings 1 through 7, the encoding sizes are set to 3, 3, 4, 5, 5, 6, and 7, respectively; for perturbation settings 1 through 7, the encoding sizes are assigned in reverse order: 7, 6, 5, 5, 4, 3, and 3.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for $10{,}000$ steps using the evaluation subset.
We assess the predictive performance of the learned representations by evaluating their ability to recover the ground-truth latent factors corresponding to their respective modalities. This evaluation accounts for the fact that some semantic variables may appear in discrete or continuous form, depending on the modality and rendering process.
Prediction performance is measured on the test set using the Matthews Correlation Coefficient (MCC) for discrete factors and the coefficient of determination ($R^{2}$) for continuous factors. Metrics are computed separately for each latent factor.
@@ E.2 Additional Results
For other latent semantic variables that are continuous in the image modality but discretized in the text modality—such as color, s_color, and b_color—we observe varying degrees of performance degradation. This drop in performance is likely attributable not only to the quantization of continuous values but also to semantic ambiguity introduced during text generation. Notably, all three attributes correspond to different aspects of color, yet they may be described using overlapping vocabulary drawn from distinct color palettes. For instance, tab:cyan from TABLEAU_COLORS refers to object color (color), cyan from CSS4_COLORS describes spotlight color (s_color), and xkcd:cyan from XKCD_COLORS indicates background color (b_color). Despite referencing different latent variables, these tokens all contain the word cyan, which is tokenized identically by nltk.PunktTokenizer, resulting in ambiguity in the text observations. These findings highlight the importance of using distinct and unambiguous content words when representing semantically different concepts in multimodal learning—particularly when the text modality is not treated merely as an auxiliary input for visual representation learning.
Interestingly, the identification of x_pos and y_pos in the text representations does not lead to improved predictability of color, in contrast to what is often observed in the image modality. This aligns with our theoretical expectation that perturbation biases disrupt the underlying causal structure in the image latent space.
@@ G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets, particularly large-scale, user-generated corpora such as LAION-5B (Schuhmann et al., 2022), often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.

### v3

@@ Results.
Figure 5: Random samples from MPI3D-Complex. Images are captured in a controlled setting.
@@ 5.3 Causal3DIdent Dataset
We further conduct experiments on the Causal3DIdent dataset (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023), a semi-synthetic dataset allowing explicitly defined causal relations among latent variables. Images are generated from 10 latent variables: 3 discrete (shape, x_pos, y_pos) and 7 continuous (color, s_pos, s_color, b_color, alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific, and the remaining as semantic variables following a causal graph in Figure 15. See Figure 7 for example images from the Causal3DIdent dataset.
@@
Training paradigm mirrors that of MPI3D-Complex. We evaluate using $R^{2}$ for continuous and MCC for discrete latent variables, averaged over three seeds. See § E.1 for full experimental details.
@@ Results.
Figure 8 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably captured across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$.
Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure.
Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in § E.2.
Figure 7: Random samples from Causal3DIdent. Images are generated from the graph in Figure 15.
@@ 7 Conclusion and Discussion
In this work, we present a formal analysis of misalignment in MMCL, examining its impact on learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, systematically discarding misaligned latent variables. When image-text pairs exhibit selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights underscore the need for multimodal learning frameworks that either mitigate misalignment or leverage beneficial biases to improve representation learning in real-world settings. A more detailed discussion of these implications is provided in App. G.
@@ C.1 Detailed Experimental Setup
We define a latent space comprising a semantic subspace $\mathcal{S}\subseteq\mathbb{R}^{10}$ and two modality-specific subspaces $\mathcal{M}_{x},\mathcal{M}_{t}\subseteq\mathbb{R}^{5}$. Variables are sampled from Gaussian priors:
$\mathbf{s}\sim\mathcal{N}(0,\Sigma_{\mathbf{s}}),\quad\mathbf{m}_{x}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{x}}),\quad\mathbf{m}_{t}\sim\mathcal{N}(0,\Sigma_{\mathbf{m}_{t}}).$
For independent semantic variables, $\Sigma_{\mathbf{s}}=I_{10}$ yields a factorized standard Gaussian, aligning with nonlinear ICA assumptions (Khemakhem et al., 2020; Sorrenson et al., 2020). For dependent semantics, we sample $\Sigma_{\mathbf{s}}$ from Wishart distribution $\mathcal{W}_{10}(I,10)$ (identity scale, 10 degrees of freedom), introducing latent dependencies as in (Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023). Modality-specific covariances $\Sigma_{\mathbf{m}_{x}}$ and $\Sigma_{\mathbf{m}_{t}}$ are sampled similarly from $\mathcal{W}_{5}(I,5)$, the Wishart distribution with identity scale and 5 degrees of freedom.
@@
Given the 10-dimensional semantic space, the total number of nonempty subsets is $10232^{10}-1=$ 1023, defining the full range of selection bias configurations $\theta$. For tractability, we choose 10 representative selection subsets $\mathbb{I}_{\theta}$, detailed in Table 2.
To isolate perturbation bias, we fix full selection ($\theta=1023$) and evaluate 10 representative perturbation subsets $\mathbb{I}_{\rho}$, as listed in Table 3. Perturbations are introduced by independently modifying each semantic variable $s_{i}$ in $\mathbb{I}_{\rho}$ with probability 0.75 using additive Gaussian noise:
$\tilde{s}_{i}=s_{i}+\mathcal{N}(0,\Sigma_{\epsilon,i}),$
where the joint noise covariance $\Sigma_{\epsilon}$ is sampled from $\mathcal{W}_{10}(I,10)$. Perturbations are applied solely to the text modality, aligning the proposed LVM and assumptions stated in § 3.
In addition to isolated biases, we consider a joint bias scenario to investigate compounding effects. Specifically, we select $\mathbb{I}_{\theta}=[8]$ (corresponding $\theta=968$, excluding the last two semantic indices), and apply perturbations to the first two indices ($\mathbb{I}_{\rho}=\{1,2\}$, correspnding to $\rho=12$).
Results across these 20 representative configurations (10 selection, 10 perturbation) and the joint bias case serve to validate the theoretical findings under a range of misalignment settings.
@@ D.1 Detailed Experimental Setup
MPI3D-Complex (Gondal et al., 2019) contains 460,800 real-world images of resolution $64\times 64\times 3$, spanning all combinations of seven mutually independent, discrete factors (see Table 6). We designate horizontal and vertical positions (hori., vert.) as image-specific due to their low-level visual nature, while treating the remaining five as semantic variables for learning and evaluation purposes.
@@
Text descriptions are generated from the ground-truth semantic attributes using content-word mappings (see Table 6). Under unbiased settings, all five semantic attributes are included. For text-specific variation, a generation template is chosen via a discrete latent variable $m_{t}\sim\mathrm{Uniform}(\{1,2,3\})$, each corresponding to a different human-written sentence structure.
@@
In selection bias settings (where the perturbable set $\mathbb{I}_{\rho}$ is empty), textual descriptions include only the content words corresponding to a subset of the true image semantic variables. We define five incremental settings: 1: {color}, 2: {color, shape}, 3: {color, shape, size}, 4: {color, shape, size, cam.}, 5: all five attributes. The specific text generation templates associated with $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ under each selection setting are detailed in Table 8.
[…]
@@ E.1 Detailed Experimental Setup
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$. For both the image and text encoders, we adopt the same architectures used in the MPI3D-Complex experiments. The encoding dimensionality is adjusted according to the bias setting: for selection settings 1 through 7, the encoding sizes are set to 3, 3, 4, 5, 5, 6, and 7, respectively; for perturbation settings 1 through 7, the encoding sizes are assigned in reverse order: 7, 6, 5, 5, 4, 3, and 3.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for $10{,}000$ steps using the evaluation subset. We assess the predictive performance of the learned representations by evaluating their ability to recover the ground-truth latent factors corresponding to their respective modalities. This evaluation accounts for the fact that some semantic variables may appear in discrete or continuous form, depending on the modality and rendering process. Prediction performance is measured on the test set using the Matthews Correlation Coefficient (MCC) for discrete factors and the coefficient of determination ($R^{2}$) for continuous factors. Metrics are computed separately for each latent factor.
@@ E.2 Additional Results
For other latent semantic variables that are continuous in the image modality but discretized in the text modality—such as color, s_color, and b_color—we observe varying degrees of performance degradation. This drop in performance is likely attributable not only to the quantization of continuous values but also to semantic ambiguity introduced during text generation. Notably, all three attributes correspond to different aspects of color, yet they may be described using overlapping vocabulary drawn from distinct color palettes. For instance, tab:cyan from TABLEAU_COLORS refers to object color (color), cyan from CSS4_COLORS describes spotlight color (s_color), and xkcd:cyan from XKCD_COLORS indicates background color (b_color). Despite referencing different latent variables, these tokens all contain the word cyan, which is tokenized identically by nltk.PunktTokenizer, resulting in ambiguity in the text observations. These findings highlight the importance of using distinct and unambiguous content words when representing semantically different concepts in multimodal learning—particularly when the text modality is not treated merely as an auxiliary input for visual representation learning. Interestingly, the identification of x_pos and y_pos in the text representations does not lead to improved predictability of color, in contrast to what is often observed in the image modality. This aligns with our theoretical expectation that perturbation biases disrupt the underlying causal structure in the image latent space.
@@ G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets—particularly large-scale, user-generated corpora such as LAION-5B (Schuhmann et al., 2022)—often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.